Zn + 2HCl ------- Zn Cl 2(subscript) +H 2subscript ?
How many moles of Zn are needed to produce 15.0 g of H(subscript 2) ? What step do you take to figure this out ?
1
Expert's answer
2020-12-21T03:56:02-0500

Molecular weight of hydrogen:

M(H2) = 1+1 = 2 g/mol

Chemical amount of hydrogen:

n(H2) = 15g/(2 g/mol) = 7.5 mol

Chemical amount of zinc:

n(Zn) = n(H2) = 7.5 mol